OpenStudy (anonymous):

Help!! Im confused on how you get x + 9 = 2(x - 1)2 into this form ax2 + bx + c = 0

OpenStudy (kirbykirby):

expand everything out. \(x+9=2(x-2)^2\\ x+9=2(x^2-4x+4)\\ x+9=2x^2-8x+8\\ 2x^2-8x-x+8-9=0\\ 2x^2-9x-1=0\)